1.Calculate the pH of a base buffer solution composed of 0.20 M NH3 and 0.15 M NH4Cl after the addition of 0.10 M HCl. The Kb of NH3 at 25oC is 1.80 × 10-5.


2.What is the pOH of a buffer solution that is 0.150 M chloroacetic acid and 0.250 M sodium chloroacetate, after the addition of 0.100 M HNO3? The value of Ka for chloroacetic acid is 1.3 x 10-3.



1
Expert's answer
2022-03-09T17:06:02-0500

1) The following reaction occurs during the addition of acid:


NH3 + HCl = NH4Cl


After the addition of acid, the new concentrations of the components will be as follows(assume that the volume of the solution remains unchanged):


С(NH3) = 0.2 - 0.1 = 0.1 M

С(NH4Cl) = 0.1 + 0.15 = 0.25 M


For base buffer solution pH = 14 + lgKb + lg(Cb/Csalt) = 8.86


2) The following reaction occurs during the addition of acid:


ClCH2COONa + HNO3 = ClCH2COOH + NaNO3


New concentrations of chloroacetic acid and sodium chloroacetate (assume that the volume of the solution remains unchanged) is:


С(ClCH2COOH) = 0.15 + 0.1 = 0.25 M

С(ClCH2COONal) = 0.25 - 0.1 = 0.15 M


For acid buffer solution pH = -l gKa - lg(Ca/Csalt) = 2.886 - 0.221 = 2.664


pOH = 14 - pH = 11.336